# On the attribution of weather events to climate change using empirically-fit extreme value distributions

### v3

@@ 3.2 Siberian Heatwave
A cumulative distribution estimate of the highest June daily-maximum temperature observed each year between 1926–2020 at the Verkhoyansk station is shown in Fig. 2a along with a 95% confidence interval of the CDF obtained by bootstrapping. In the first step of our analysis, we consider cumulative distribution function fits to the observed distribution, assuming the data follows either a GEV or normal distribution. The fitted distributions are specified here to be stationary (by setting $\alpha=0$ and $\mu=\mu_{0}$ in Eqns. 14). Fits are obtained by maximizing the likelihood of the data, and the results are shown in Fig. 2a. Both distributions generally fit the observations well, with the normal fit completely within the 95% range of the observed distribution, and with the GEV fit being good except at the highest temperature, where it slightly underestimates the probability of obtaining the highest temperatures. There is only one data point above 35 °C, so the empirical CDF is poorly constrained there in any case. Interestingly, the normal distribution and GEV CDFs look similar and lie within the confidence intervals defined for the empirical CDF, despite the normal distribution having fewer parameters than the GEV. It is worth noting that a GEV and a normal distribution may assign very different probabilities to extreme events, even when the two distributions provide an adequate fit to the observed CDF. These modeling decisions are therefore of critical importance for attribution studies, which usually focus as much or more on estimation of the exceedance probability and return period of an event — and, consequently, on the changing likelihood and fraction of attributable risk of an event — as they do on the change in intensity. It is therefore useful to consider alternate distributional forms and, in many ways, the normal is the simplest and often a useful default.
@@
The quality of fits indicates that an extreme value distribution is not empirically a better choice than a normal distribution for fitting the data. Similarly, a non-stationary distribution with $\alpha\neq 0$ may not be necessary given that the stationary distributions seem to provide a satisfactory fit to the empirical cumulative distributions, and we further examine this next. CDFs are commonly used to differentiate between distribution functions, and the results of a Kolmogorov-Smirnov test (Supplementary Fig. SI-1 for normal, GEV and empirical distributions) also suggests that the normal distribution is consistent with the observed data. An application of Occam’s razor principle suggests that the simpler normal distribution should be used rather than a GEV if there is not a significant improvement in the distributional fit with the GEV. This is also discussed by Philip et al. (2020) who mention explicitly that the extreme value distributions may not outperform a simpler (normal) distribution in some cases. Otto et al. (2018) opted to use a normal distribution in lieu of a GPD for attribution analysis of a cold spell in Peru because the normal “uses all data, rather than the GPD, which samples only the cold tail”. CC2021 found that a normal distribution fits the data well but proceeded to use a GEV for the bulk of the analysis. These results question the need for extreme value distribution functions when it comes to attribution analysis. One might want to use proper selection criteria (the Akaike Information Criterion, AIC, or deviance statistic or similar) to rigorously decide which of the models performs best.
@@
The fact that the LE results indicate a stronger link between Siberian extremes and GMST (that is, a larger $\alpha$) is consistent with the LE, on average, simulating an Arctic-wide summer warming rate of 1.8 C per C warming in GMST, whereas an observational analysis only indicates a rate of 1.3 C Arctic JJA warming per C in GMST (using GISTEMP for both; Fig. 4a,b). This greater amplification of Arctic warming may be related to the fact that CESM1 has a relatively high climate sensitivity of 4.1 K Gettelman et al. (2019). The WWA approach leverages a range of climate models so that results are not dependent entirely on one GCM. That said, one expects heatwaves to be strongly influenced by the JJA climatology, and Fig. 4c shows that different ensemble members, corresponding to different realizations of internal variability, show very different dependencies of the Arctic JJA climatology on GMST. It is difficult, therefore, to estimate from the single realization corresponding to the observed record to what degree the Arctic JJA climatology is affected by ACC versus internal variability, adding to the uncertainty in attributing Siberian heatwaves. To address that, the WWA protocol includes a step where the same analysis is applied to climate models, and results are reported from both sources.
@@
Figure 4: Internal variability and the attribution of Siberian heatwaves. (a) A scatter plot of the annual average JJA Arctic (north of $70^{\circ}$ N) surface temperatures against the 4-year running mean of GISTEMP annual GMST. (b) Same, for the 35 members of the CESM Large Ensemble over the period 1926–2019. (c) PDF of the slopes of annual average JJA Arctic surface temperatures against the 4-year running mean of annual GMST within each member of the CESM Large Ensemble with 5,000 bootstrap resamples (red). The same distribution is indicated for GISTEMP observations in blue. The vertical red line indicates the average slope across the CESM Large Ensemble, and the vertical blue line indicates the slope determined from observations.
@@ 3.3 Australian Bushfire Heatwave
Figure 7: Australian Bushfire Heatwave: distribution of $\alpha$ values. (a) Scatter plot of the highest annual 7-day running mean of daily maximum temperatures against GMST (following van Oldenborgh et al., 2021). The red line indicates a linear regression fit to the data, and the blue line indicates a linear fit using the location parameter and $\alpha$ derived from the GEV maximum likelihood estimate. (b; red) PDF of the GEV-fit $\alpha$ distributions from observations, using bootstrap resampling. The $\alpha$ value from the observations without resampling is shown by the red line (at 1.7) (b; blue) Effect of the internal variability and number of data from a model large ensemble on the uncertainty in $\alpha$ for the Australian Bushfire heatwave example. Shown is a histogram of the GEV-fit $\alpha$ PDFs derived from the $\alpha$ value from each ensemble member with no resampling. The average $\alpha$ value across the Large Ensemble is indicated with the blue line.
@@ 4 Conclusions
We were inspired in this work by a series of attribution studies by the “World Weather Attribution Project” (WWA) and evaluated the part of their methodology that involves a fit of extreme value distribution functions to observations. The full protocol used by this effort Philip et al. (2020) involves additional tools and model analyses, including selecting and defining extreme events, analyzing trends in observational data, using multi-model ensembles to simulate event scenarios with and without human-induced climate change, performing quality checks on model accuracy, and synthesizing results to determine climate change’s influence on the event’s intensity and likelihood. A central WWA method nevertheless involves fitting an extreme value distribution function to an observed record, with the distribution parameters depending on global mean surface temperature (GMST), whose variation is taken to represent ACC. For example, the location parameter of an extreme value distribution (GEV) is sometimes made a linear function of GMST, $\overline{T}$, as $\mu=\mu_{0}+\alpha\times\overline{T}$. The GMST dependence included in the function is then used to calculate the change in return time of extreme events. This method was applied to many extreme events such as heatwaves, cold spells, floods, droughts, and more (e.g., van Oldenborgh et al., 2016; van der Wiel et al., 2017; van Oldenborgh et al., 2017; Ciavarella et al., 2021; van Oldenborgh et al., 2021; Harrington et al., 2022). We examined the robustness of the results of three example WWA studies, demonstrated when the results may be unreliable due to the effects of internal variability, and suggested several additional statistical tests, reviewed below, that may increase confidence in the results of such an analysis.
@@
We began by examining the GMST dependence of extreme heatwave events in Siberia (following CC2021) and Australia (following OK2021), as well as drought in Madagascar (following OK2021), using data from a CESM preindustrial simulation. We found a significant GMST dependence in the first two cases, indicating that such dependence can result from internal variability rather than only represent the effects of ACC. The GMST dependence, in these cases, is a result of regional internal variability such as ENSO or other slowly varying modes, which affects the extreme events being examined, and also weakly affects GMST. This points that the value of $\alpha$ calculated in the more realistic case where both effects are present includes the effects of internal variability in addition to those of ACC. The value of $\alpha$ in the Australian case is especially large, and could naively be interpreted as indicating that every degree increase in GMST is associated with a 5-degree increase in Australian heatwaves. Heuristically, the large values of $\alpha$ can be thought of as a result of internal variability (possibly ENSO) affecting GMST by 0.2 degrees for every 1-degree effect it has on Australian heatwaves. As noted earlier, this issue could be mitigated in part by including indices for modes of internal variability as covariates in the model fitting.
@@
An important issue raised by this work is the discrepancy between the modeled and observed findings, as seen, for example, in the deviance statistic results for the observations versus CESM Large Ensemble (Figs. 2b, 6b, 8b). This may be due to a particular sequence of internal variability in the observations, measurement error in the observational record, or issues with the simulation of extreme weather in the CESM Large Ensemble. It would be helpful to test if these discrepancies also occur with other models. The use of multiple models by the WWA (Philip et al. (2020), OK2021), may help mitigate against this kind of problem.
Overall, our results suggest uncertainty in the interpretation of the dependence of the parameters of extreme event distributions on the global mean surface temperature as reflecting ACC. The proposed and demonstrated deviance statistic, MRL plot, and a careful analysis of uncertainty ranges, especially for $\alpha$, may help avoid misguided confidence in attribution results. As noted in the introduction, WWA does not exclusively rely upon the empirical methodology evaluated here in order to reach conclusions, and also evaluates other historical data, model simulations, and the dynamical context. Philip et al. (2020) also highlight some caveats associated with their study, including the necessity of testing for other probability laws (e.g., Gaussian, Gamma, etc.) and covariates (ENSO, PDO, etc.), among other issues. Nevertheless, the observations-based component of the WWA methodology examined here still figures prominently in their analyses, and we have highlighted several caveats with its interpretation, foremost that internal variability can give rise to the appearance of strong covariance between GMST and local extremes. It is important to study when the ACC signal emerges distinctly relative to internal variability on a case-by-case basis. The ACC signal is known to have emerged in most regions of the world for temperature extremes, while this may not be the case for variables like precipitation that have much higher variability. Time of emergence is a factor that should always be considered when discussing the results of an attribution study. Our results do not provide evidence against a relationship between GMST and local extremes but, rather, that internal variability needs to be taken into account when fitting a distribution function to the record. As currently implemented, the WWA approach can lead to first-order biases in some cases. It, therefore, seems that it may be best to use a multi-method approach to ensure a robust attribution approach, including the storyline approach that can add insight about the change to extreme events (e.g., Shepherd et al., 2018; Leach et al., 2024; Robin and Ribes, 2020b; Sippel et al., 2024).

### v4

@@ 3.2 Siberian Heatwave
A cumulative distribution estimate of the highest June daily-maximum temperature observed each year between 1926–2020 at the Verkhoyansk station is shown in Fig. 2a along with a 95% confidence interval of the CDF obtained by bootstrapping. In the first step of our analysis, we consider cumulative distribution function fits to the observed distribution, assuming the data follows either a GEV or normal distribution. The fitted distributions are specified here to be stationary (by setting $\alpha=0$ and $\mu=\mu_{0}$ in Eqns. 14). Fits are obtained by maximizing the likelihood of the data, and the results are shown in Fig. 2a. Both distributions generally fit the observations well, with the normal fit completely within the 95% range of the observed distribution, and with the GEV fit being good except at the highest temperature, where it slightly underestimates the probability of obtaining the highest temperatures. There is only one data point above 35 °C, so the empirical CDF is poorly constrained there in any case. Interestingly, the normal distribution and GEV CDFs look similar and lie within the confidence intervals defined for the empirical CDF, despite the normal distribution having fewer parameters than the GEV. It is worth noting that a GEV and a normal distribution may assign very different probabilities to extreme events, even when the two distributions provide an adequate fit to the observed CDF. These modeling decisions are therefore of critical importance for attribution studies, which usually focus as much or more on estimation of the exceedance probability and return period of an event — and, consequently, on the changing likelihood and fraction of attributable risk of an event — as they do on the change in intensity. It is therefore useful to consider alternate distributional forms.
@@
The quality of fits indicates that an extreme value distribution is not empirically a better choice than a normal distribution for fitting the data. Similarly, a non-stationary distribution with $\alpha\neq 0$ may not be necessary given that the stationary distributions seem to provide a satisfactory fit to the empirical cumulative distributions, and we further examine this next. CDFs are commonly used to differentiate between distribution functions, and the results of a Kolmogorov-Smirnov test (Supplementary Fig. SI-1 for normal, GEV and empirical distributions) also suggests that the normal distribution is consistent with the observed data. An application of Occam’s razor principle suggests that the simpler normal distribution should be used rather than a GEV if there is not a significant improvement in the distributional fit with the GEV. This is also discussed by Philip et al. (2020) who mention explicitly that the extreme value distributions may not outperform a simpler (normal) distribution in some cases. Otto et al. (2018) opted to use a normal distribution in lieu of a GPD for attribution analysis of a cold spell in Peru because the normal “uses all data, rather than the GPD, which samples only the cold tail”. CC2021 found that a normal distribution fits the data well but proceeded to use a GEV for the bulk of the analysis. These results question the need for extreme value distribution functions when it comes to attribution analysis. One might want to use proper selection criteria (the Akaike Information Criterion, AIC, or the deviance statistic AKA the log likelihood ratio test which we use in this work, or similar) to rigorously decide which of the models performs best.
@@
The fact that the LE results indicate a stronger link between Siberian extremes and GMST (that is, a larger $\alpha$) is consistent with the LE, on average, simulating an Arctic-wide summer warming rate of 1.8 C per C warming in GMST, whereas an observational analysis only indicates a rate of 1.3 C Arctic JJA warming per C in GMST (using GISTEMP for both; Fig. 4a,b). This greater amplification of Arctic warming may be related to the fact that CESM1 has a relatively high climate sensitivity of 4.1 K Gettelman et al. (2019). That said, one expects heatwaves to be strongly influenced by the JJA climatology, and Fig. 4c shows that different ensemble members, corresponding to different realizations of internal variability, show very different dependencies of the Arctic JJA climatology on GMST. It is difficult, therefore, to estimate from the single realization corresponding to the observed record to what degree the Arctic JJA climatology is affected by ACC versus internal variability, adding to the uncertainty in attributing Siberian heatwaves. In order to address this issue, the WWA protocol includes a step where the same analysis is applied to climate models, leveraging a range of climate models so that results are not dependent entirely on one GCM, and results are reported from both sources.
@@
Figure 4: Internal variability and the attribution of Siberian heatwaves. (a) A scatter plot of the annual average JJA Arctic (north of $70^{\circ}$ N) surface temperatures against the 4-year running mean of GISTEMP annual GMST. (b) Same, for the 35 members of the CESM Large Ensemble over the period 1926–2019. (c) PDF of the slopes of annual average JJA Arctic surface temperatures against the 4-year running mean of annual GMST within each member of the CESM Large Ensemble with 5,000 bootstrap resamples (red). The same distribution is indicated for GISTEMP observations in blue. The vertical red line indicates the mean slope across the CESM Large Ensemble, and the vertical blue line indicates the slope determined from observations.
@@ 3.3 Australian Bushfire Heatwave
Figure 7: Australian Bushfire Heatwave: distribution of $\alpha$ values. (a) Scatter plot of the highest annual 7-day running mean of daily maximum temperatures against GMST (following van Oldenborgh et al., 2021). The red line indicates a linear regression fit to the data, and the blue line indicates a linear fit using the location parameter and $\alpha$ derived from the GEV maximum likelihood estimate. (b; red) PDF of the GEV-fit $\alpha$ distributions from observations, using bootstrap resampling. The $\alpha$ value from the observations without resampling is shown by the red line (at 1.7) (b; blue) Effect of the internal variability and number of data from a model large ensemble on the uncertainty in $\alpha$ for the Australian Bushfire heatwave example. Shown is a histogram of the GEV-fit $\alpha$ PDFs derived from the $\alpha$ value from each ensemble member with no resampling. The mean $\alpha$ value across the Large Ensemble is indicated with the blue line.
@@ 4 Conclusions
We were inspired in this work by a series of attribution studies by the “World Weather Attribution initiative” (WWA) and evaluated the part of their methodology that involves a fit of extreme value distribution functions to observations. The full protocol used by this effort Philip et al. (2020) involves additional tools and model analyses, including selecting and defining extreme events, analyzing trends in observational data, using multi-model ensembles to simulate event scenarios with and without human-induced climate change, performing quality checks on model accuracy, and synthesizing results to determine climate change’s influence on the event’s intensity and likelihood. A central WWA method nevertheless involves fitting an extreme value distribution function to an observed record, with the distribution parameters depending on global mean surface temperature (GMST), whose variation is taken to represent ACC. For example, the location parameter of an extreme value distribution (GEV) is sometimes made a linear function of GMST, $\overline{T}$, as $\mu=\mu_{0}+\alpha\times\overline{T}$. The GMST dependence included in the function is then used to calculate the change in return time of extreme events. This method was applied to many extreme events such as heatwaves, cold spells, floods, droughts, and more (e.g., van Oldenborgh et al., 2016; van der Wiel et al., 2017; van Oldenborgh et al., 2017; Ciavarella et al., 2021; van Oldenborgh et al., 2021; Harrington et al., 2022). We examined the robustness of the results of three example WWA studies, demonstrated when the results may be unreliable due to the effects of internal variability, and suggested several additional statistical tests, reviewed below, that may increase confidence in the results of such an analysis.
@@
We began by examining the GMST dependence of extreme heatwave events in Siberia (following CC2021) and Australia (following OK2021), as well as drought in Madagascar (following OK2021), using data from a CESM preindustrial simulation. We found a significant GMST dependence in the first two cases, indicating that such dependence can result from internal variability rather than only represent the effects of ACC. The GMST dependence, in these cases, is a result of regional internal variability such as ENSO or other climate variability modes, which affects the extreme events being examined, and also weakly affects GMST. This points that the value of $\alpha$ calculated in the more realistic case where both effects are present includes the effects of internal variability in addition to those of ACC. The value of $\alpha$ in the Australian case is especially large, and could naively be interpreted as indicating that every degree increase in GMST is associated with a 5-degree increase in Australian heatwaves. Heuristically, the large values of $\alpha$ can be thought of as a result of internal variability (possibly ENSO) affecting GMST by 0.2 degrees for every 1-degree effect it has on Australian heatwaves. As noted earlier, this issue could be mitigated in part by including indices for modes of internal variability as covariates in the model fitting.
@@
An important issue raised by this work is the discrepancy between the modeled and observed findings, as seen, for example, in the deviance statistic results for the observations versus CESM Large Ensemble (Figs. 2b, 6b, 8b). This may be due to a particular sequence of internal variability in the observations, measurement error in the observational record, or issues with the simulation of extreme weather in the CESM Large Ensemble. It would be helpful to test if these discrepancies also occur with other models. The use of multiple models by WWA (Philip et al. (2020), OK2021), may help mitigate against this kind of problem.
Overall, our results suggest uncertainty in the interpretation of the dependence of the parameters of extreme event distributions on the global mean surface temperature as reflecting ACC. The proposed and demonstrated deviance statistic, MRL plot, and a careful analysis of uncertainty ranges, especially for $\alpha$, may help avoid misguided confidence in attribution results. As noted in the introduction, WWA does not exclusively rely upon the empirical methodology evaluated here in order to reach conclusions, and also evaluates other historical data, model simulations, and the dynamical context. Philip et al. (2020) also highlight some caveats, including the necessity of testing for other probability laws (e.g., Gaussian, Gamma, etc.) and covariates (ENSO, PDO, etc.), among other issues. Nevertheless, the observations-based component of the WWA methodology examined here still figures prominently in their analyses, and we have highlighted several caveats with its interpretation, foremost that internal variability can give rise to the appearance of strong covariance between GMST and local extremes. It is important to study when the ACC signal emerges distinctly relative to internal variability on a case-by-case basis. The ACC signal is known to have emerged in most regions of the world for temperature extremes, while this may not be the case for variables like precipitation that have much higher variability. Time of emergence is a factor that should always be considered when discussing the results of an attribution study. Our results do not provide evidence against a relationship between GMST and local extremes but, rather, that internal variability needs to be taken into account when fitting a distribution function to the record. As currently implemented, the WWA approach can lead to first-order biases in some cases. It, therefore, seems that it may be best to use a multi-method approach to ensure a robust attribution approach (e.g., Shepherd et al., 2018; Leach et al., 2024; Robin and Ribes, 2020b; Sippel et al., 2024).